OpenStudy (anonymous):

Y>x^2 +2x -4 Y>x^2 -4x -1 (These are parabolas!) i need to solve, graph, and shade the parabolas! Please help!

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

to graph y >x^2 +2x -4, you need to graph y = x^2 +2x -4 first

jimthompson5910 (jim_thompson5910):

to graph y = x^2 + 2x - 4, you plot a series of points to get a rough idea of what it's shape is, then you draw a curve (your best guess) through them all

OpenStudy (anonymous):

How do i get the points?

jimthompson5910 (jim_thompson5910):

To get any point, you plug in a number for x to get corresponding value for y. For example, let's say x = 0 y = x^2 + 2x - 4 y = (0)^2 + 2(0) - 4 ... replace x with 0 (since x = 0) y = 0 + 2(0) - 4 y = 0 + 0 - 4 y = -4 So when x = 0, y = -4 giving us the point (0,-4) So that means the point (0,-4) is on the graph of y = x^2 + 2x - 4

jimthompson5910 (jim_thompson5910):

To get other points, you use other values for x (say x = 1, x = 2, etc)

OpenStudy (anonymous):

Is threesome formula for finding the vertex?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Is there*

jimthompson5910 (jim_thompson5910):

x = -b/(2a) is the axis of symmetry formula

OpenStudy (anonymous):

Ok, so use that formula to be the vertex,

jimthompson5910 (jim_thompson5910):

yes, the axis of symmetry is the x coordinate of the vertex

OpenStudy (anonymous):

Then i use other numbers to plug into the equation?

jimthompson5910 (jim_thompson5910):

once you know the x coordinate of the vertex, you plug it back into the original equation to find the y coordinate of the vertex

OpenStudy (anonymous):

And i have to shade

OpenStudy (anonymous):

Where we pic a test point

jimthompson5910 (jim_thompson5910):

after you graph the parabola for y = x^2 + 2x - 4 you shade based on what the inequality sign is

jimthompson5910 (jim_thompson5910):

In the case of Y>x^2 +2x -4, we shade above the curve because we have a greater than sign

jimthompson5910 (jim_thompson5910):

if the sign was <, you'd shade below the curve

OpenStudy (anonymous):

But what if the y is <

jimthompson5910 (jim_thompson5910):

then you shade below

OpenStudy (anonymous):

Imconfuseh

jimthompson5910 (jim_thompson5910):

you could use the test point method, but this way is faster

OpenStudy (anonymous):

Actually i goit thanks

jimthompson5910 (jim_thompson5910):

yw